# Nash Equilibria in Games with Playerwise Concave Coupling Constraints: Existence and Computation

### v1

@@ 4 LEARNING CONSTRAINED NASH EQUILIBRIA
In this section, we turn to the problem of finding approximate constrained Nash equilibria in a deterministic, gradient-based, independent learning setting.
First, we introduce an approximate notion of constrained Nash equilibrium that serves as the solution concept for our algorithms. Note that the case $\epsilon=0$ recovers the exact notion from Section 2.
@@
For $\epsilon>0$, $x\in\mathcal{X}$ is a constrained $\epsilon$-approximate Nash equilibrium if $x\in\mathcal{C}$ and for all $i\in[m]$ and $x^{\prime}_{i}\in\mathcal{C}_{i}(x_{-i})$, it holds that $u_{i}(x^{\prime}_{i},x_{-i})\leq u_{i}(x)+\epsilon$.
While Assumptions 1 and 2 suffice for existence, without additional assumptions on the game structure, the problem of finding $\epsilon$-approximate Nash equilibria is known to be PPAD-hard even for unconstrained normal-form games (Daskalakis et al.,, 2009; Rubinstein,, 2015). Therefore, in Section 4.1, we focus on the commonly studied tractable sublass of potential games. Before specializing to this case, we introduce the considered independent learning protocol, discuss challenges and additional assumptions, and outline our general approach.
Independent learning protocol
We assume players to interact with the game for $T$ rounds. In each round $t$, for $0\leq t\leq T-1$, player $i$ chooses some strategy $\smash{x_{i}^{(t)}\in\mathcal{X}^{i}}$ and receives first-order feedback for the utility, i.e. $\nabla_{x_{i}}u_{i}(x^{(t)})$, as well as zeroth- and first-order feedback for each constraint $j\in[b]$, i.e., $c_{j}(x^{(t)})$ and $\nabla_{x_{i}}c_{j}(x^{(t)})$. Players do not observe other players’ strategies or utility feedback. Moreover, besides the game interaction, no communication or coordination (e.g., shared randomness) among players is allowed.
Smoothness and Lipschitz continuity of utility functions are standard assumptions in gradient-based learning, see e.g. Mertikopoulos et al., (2024); we extend these regularity conditions to the constraints.
@@
There exist constants $L$ and $M$ such that the functions $u_{i}$ for $i\in[m]$ and $c_{j}$ for $j\in[b]$ are all differentiable, $L$-Lipschitz, and $M$-smooth in the joint strategies $x\in\mathcal{X}$.
Moreover, we make the following assumption on strict feasibility of the initial strategy.
Assumption 4.
There exist $\gamma>0$ and $x^{(0)}\in\mathcal{X}$ such that $\min_{j\in[b]}\{c_{j}(x^{(0)})-\alpha_{j}\}\geq\gamma$. For all $i\in[m]$, we assume $x^{(0)}_{i}$ is initially known to player $i$.
Note that without assuming initial feasibility, finding a feasible strategy in a two-player game with bilinear constraints is as hard as finding a global maximizer of a bilinear function. The latter problem is known to be NP-hard  (Witsenhausen,, 1986) even in a centralized setting.
Furthermore, preserving the initial feasibility across subsequent iterates motivates the use of interior point methods. Our method is inspired by log barrier regularized gradient descent, an interior point method that has recently been studied in the context of nonconvex optimization under nonconvex constraints (Hinder and Ye,, 2024; Usmanova et al.,, 2024).
@@
where $\eta>0$ is a parameter controlling the regularization. Observe that for each constraint $j$, the log-term serves as a barrier that tends to -$\infty$ as $c_{j}(z)$ approaches the threshold $\alpha_{j}$ for any $j\in[b]$. Then, we propose the following independent log barrier regularized gradient ascent: Suppose in iteration $t$ for $0\leq t\leq T-1$, each player $i$ simultaneously performs the projected regularized gradient update
@@
Smoothness along trajectory
Even though we assume smoothness of utilities and constraints, $B^{\eta}_{i}(x)$ is not globally smooth due to unbounded growth of the log-term towards the boundary of $\mathcal{C}$. However, we show that $B^{\eta}_{i}(x)$ is smooth over the joint strategy trajectory of independent log barrier regularized gradient ascent when using suitable adaptive stepsizes $\smash{\gamma^{(t)}}$.
@@
For $\rho>0$, let $\mathcal{J}_{\rho}(x)\coloneqq\{j\in[b]\mid c_{j}(x)-\alpha_{j}\leq\rho\}$ be the set of $\rho$-approximately active constraints for strategy $x\in\mathcal{C}$. For some $\rho\in(0,\frac{\gamma}{2}]$, there exists $\ell>0$, such that for all $x\in\mathcal{X}$ and all $j\in\mathcal{J}_{\rho}(x)$, there exists a direction $\delta$ with $x+\delta\in\mathcal{X}$, $\lVert\delta\rVert\leq 1$, such that $\langle\delta,\nabla c_{j}(x)\rangle>\ell$.
@@
4.1 Learning in Constrained Potential Games
We consider the class of games with coupling constraints that, when the constraints are omitted, reduce to a potential game, as introduced by Monderer and Shapley, (1996). For our purpose, it suffices to have a potential function defined over the feasible region.
Assumption 6.
Suppose there exists a potential function $\Phi:\mathcal{C}\to\mathbb{R}$ such that for all players $i\in[m]$, strategy profiles $x\in\mathcal{C}$, and playerwise deviations $x^{\prime}_{i}\in\mathcal{C}_{i}(x_{-i})$,
$\Phi(x)-\Phi(x^{\prime}_{i},x_{-i})=u_{i}(x)-u_{i}(x^{\prime}_{i},x_{-i}).$
Examples of such continuous potential games trivially include identical interest games, and, as shown in Example 3, mixed strategy extensions of finite action potential games. This class of potential games also arises in economics, e.g. in the context of the Cournot competition (Cournot,, 1838), in communication networks (Scutari et al.,, 2006), and in transportation (Altman and Wynter,, 2004).
Example 3.
Consider the unconstrained game with joint action space $\mathcal{A}\coloneqq\times_{i\in[m]}\mathcal{A}_{i}$ where each $\mathcal{A}_{i}$ is a finite set, and utilities $\widetilde{u}_{i}:\mathcal{A}_{i}\to[0,1]$ such that there exists a potential function $\phi:\mathcal{A}\to\mathbb{R}$ that satisfies $\phi(a)-\phi(a_{i}^{\prime},a_{-i})=\widetilde{u}_{i}(a)-\widetilde{u}_{i}(a_{i}^{\prime},a_{-i})$ for all $i\in[m]$, $a\in\mathcal{A}$, and $a^{\prime}_{i}\in\mathcal{A}_{i}$. Then, the mixed strategy extension of this game with strategy space $\mathcal{X}\coloneqq\times_{i\in[m]}\Delta(\mathcal{A}_{i})$ and utilities $u_{i}(x)\coloneqq\mathbb{E}_{a\sim x}[\widetilde{u}_{i}(a)]$ is a potential game with potential $\Phi(x)\coloneqq\mathbb{E}_{a\sim x}[\phi(a)]$. See Appendix F for a proof of this fact.
Next, we state our best-iterate convergence result for potential games with coupling constraints. An overview of the analysis is given below; for the full proof, we refer to Appendix F. {restatable}theoremthmpotential Let $\epsilon>0$ and set $\eta=\epsilon$. Let Assumptions 1, 3, 4, 6, and 5 hold with $\rho\geq\eta$. For stepsizes $\{\gamma^{(t)}\}_{t=0}^{T-1}$ chosen as in (2), suppose we perform $T=\mathcal{O}(\epsilon^{-3})$ rounds of simultaneous updates as in (1). Then, there exists $0\leq t\leq T$ such that $x^{(t)}$ is a constrained $\epsilon$-approximate Nash equilibrium.
Remark 2.
@@ 4.1 Learning in Constrained Potential Games
    Compared to the corresponding result in unconstrained potential games, see Theorem B.6 of Anagnostides et al., (2022) which achieves a $\mathcal{O}(\epsilon^{-2})$ best-iterate rate, log barrier regularization comes at the cost of an additional $\mathcal{O}(\epsilon^{-1})$ factor. This is due to small stepsizes as iterates approach the boundary of the feasible region.
@@
    While we are not aware of any results on learning static potential games with playerwise concave coupling constraints, a related result exists for constrained Markov potential games under an independent learning protocol similar to ours (Jordan et al.,, 2024). Unlike our approach, it relies on a double-loop proximal-point-like algorithm and, in the exact-gradient setting, achieves best-iterate convergence to a constrained $\epsilon$-approximate Nash equilibrium a slower $\mathcal{O}(\epsilon^{-4})$ rate. Moreover, Jordan et al., (2024) requires a so-called uniform Slater’s condition which in practice may be harder to verify than our Assumption 5.
@@
First, we observe that due to the identical interest structure of the log barrier regularization, $\Phi^{\eta}(x)\coloneqq\Phi(x)+\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$ is a potential function for the regularized game. Therefore, partial gradients are aligned with the potential gradient, i.e., $\nabla_{x_{i}}B^{\eta}_{i}(x)=\nabla_{x_{i}}\Phi^{\eta}(x)$ for all $i\in[m]$ which allows us to equivalently analyze our independent updates as a centralized log barrier regularized gradient ascent on $\Phi^{\eta}$.
Next, we argue about smoothness of $\Phi^{\eta}$ along the gradient ascent trajectory. We say $x\in\mathcal{C}$ is $\beta$-feasible if $c_{j}(x)-\alpha_{j}\geq\beta$ for all $j\in[b]$. For $\beta$-feasible $x$, we give a Hessian bound $\lVert\nabla^{2}\Phi^{\eta}(x)\rVert\leq M^{\eta,\beta}_{\Phi}$ where $M^{\eta,\beta}_{\Phi}\coloneqq M\sqrt{m}+\eta Mb\beta^{-1}+\eta L^{2}b\beta^{-2}$. Moreover, we show that if an iterate $x^{(t)}$ is $\beta^{(t)}$-feasible, and the stepsize sequence is chosen adaptively such that for all $0\leq t\leq T-1$, it holds that $0<\gamma^{(t)}\leq\min_{j\in[b]}\frac{c_{j}(x^{(t)})-\alpha_{j}}{2mL^{2}}$, then all strategies on the line segment $x^{(t)},x^{(t+1)}]$ are $\beta^{(t)}/2$-feasible. Consequently, $\Phi^{\eta}$ is $M_{\Phi}^{\eta,\beta^{(t)}/2}$-smooth over $x^{(t)},x^{(t+1)}]$. Therefore, if we choose stepsizes
a standard inequality from nonconvex smooth optimization (Bubeck et al., (2015), Lemma 3.6) shows sufficient increase proportional to $\gamma^{(t)}$. Nevertheless, $M_{\Phi}^{\eta,\beta^{(t)}/2}$ may grow arbitrarily and lead to diminishing stepsizes. Leveraging Assumption 5, we give a lower bound on $\gamma^{(t)}$ based on the fact that the log barrier term dominates and steers iterates away from the boundary of $\mathcal{C}$ whenever the feasibility margin $\beta$ is too small.
As a result, for our choices of $\eta$ and $T$, we obtain best-iterate convergence at an $\mathcal{O}(\epsilon^{-3})$-rate towards an $\epsilon$-stationary point $x^{\star}$ of $\Phi^{\eta}$. We show that this implies a set of $\epsilon$-approximate KKT conditions to hold for the constrained optimization problem $\max_{x\in\mathcal{C}}\Phi^{\eta}(x)$ at $x^{\star}$, which in turn implies $x^{\star}$ to be an $\epsilon$-approximate constrained Nash equilibrium.
@@ 5 SIMULATIONS
To demonstrate the convergence of our method, we conduct numerical simulations on two continuous potential games with shared coupling constraints. The first is a simple two-player identical interest game played over a nonconvex feasible region. The second is a routing problem with link capacity constraints.
Identical interest game
@@
Figure 3: Strategy trajectory in the utility landscape of the identical interest game; colors indicate the log barrier value; grey is infeasible.
Starting from an initial feasible point, we apply our log barrier regularized independent gradient ascent to the constrained potential game. Figure 3 visualizes the joint strategy trajectory within the utility landscape where colors indicate the log barrier value. The trajectory (in red) remains within the nonconvex feasible region throughout the iterations, aligning with our theoretical observations. In Figure 4, we plot the $\text{Nash-Gap}_{i}(x)\coloneqq\max_{x_{i}^{\prime}\in\mathcal{C}_{i}(x_{-i})}u_{i}(x_{i}^{\prime},x_{-i})-u_{i}(x)$ for each player $i$. This can be computed by solving the single-player constrained optimization problem where $x_{-i}$ remains fixed. The Nash gap serves as a measure of convergence, as for any $\epsilon>0$, if $\text{Nash-Gap}(x)\coloneqq\max_{i\in[m]}\text{Nash-Gap}_{i}(x)\leq\epsilon$, then $x$ is a constrained $\epsilon$-approximate Nash equilibrium. Hence, we observe convergence to a strategy near the boundary of $\mathcal{C}$ that forms an approximate constrained Nash equilibrium.
Figure 4: Nash gaps and constraint values over iterations $t$ for the identical interest game.
@@
Moving beyond the identical interest case, we build upon an unconstrained distributed routing problem that has been shown to satisfy the potential structure (see Example 2.2, Tatarenko, (2017)). We extend this model by incorporating coupled capacity constraints. Specifically, 5 players, one corresponding to each color in Figure 5 aim to route some amount of flow through 5 paths in the network, each associated to the respective player ($R_{1}$: blue, $R_{2}$: red, $R_{3}$: brown, $R_{4}$: orange, $R_{5}$: green). For $i\in[5]$, we let $x_{i}\in[0,10]$ denote the flow on player $i$’s path $R_{i}$.
@@
Starting from feasible flows $x_{i}=\frac{1}{2}$ for all $i\in[5]$, Figure 6 shows the convergence of log barrier regularized independent gradient ascent to a constrained Nash equilibrium of the routing game. Note that here $\text{Nash-Gap}(x^{(t)})\coloneqq\max_{i\in[5]}\text{Nash-Gap}_{i}(x^{(t)})$. We point out that, again, all iterates remain within the feasible region.

### v2

@@ 4 LEARNING CONSTRAINED NASH EQUILIBRIA
If players update their strategies using only local, gradient-based information, can they efficiently learn an approximate constrained Nash equilibrium? In this section, we investigate this question in potential games with shared coupling constraints, and propose methods for finding approximate constrained Nash equilibria in a deterministic, independent learning setting.
4.1 Constrained potential games
While Assumptions 1 and 2 suffice for existence, without additional assumptions on the game structure, the problem of finding $\epsilon$-approximate Nash equilibria is known to be PPAD-hard even for unconstrained normal-form games (Daskalakis et al.,, 2009; Rubinstein,, 2015). Therefore, we focus on the subclass of potential games (Monderer and Shapley,, 1996) with coupling constraints. For our purpose, it suffices to have a potential function defined over the feasible region.
@@ 4.1 Constrained potential games
A potential game with coupling constraints is a constrained game for which there exists a function $\Phi:\mathcal{C}\to\mathbb{R}$ such that for all players $i\in[m]$, strategy profiles $x\in\mathcal{C}$, and $x^{\prime}_{i}\in\mathcal{C}_{i}(x_{-i})$, we have
$\Phi(x)-\Phi(x^{\prime}_{i},x_{-i})=u_{i}(x)-u_{i}(x^{\prime}_{i},x_{-i}).$
Examples of such continuous potential games trivially include cooperative games, and, as shown in Example 3 below, mixed strategy extensions of finite action potential games with coupling constraints. This class of potential games also arises in economics, e.g. in the context of the Cournot competition (Cournot,, 1838), in communication networks (Scutari et al.,, 2006), and in transportation (Altman and Wynter,, 2004).
Example 3.
Consider the game with joint action space $\mathcal{A}\coloneqq\times_{i\in[m]}\mathcal{A}_{i}$ where each $\mathcal{A}_{i}$ is a finite set, utilities $\widetilde{u}_{i}:\mathcal{A}\to[0,1]$ for $i\in[m]$, and constraints $\widetilde{c}_{j}:\mathcal{A}\to[0,1]$ with thresholds $\alpha_{j}$, for $j\in[b]$, that induce a feasible region $\widetilde{\mathcal{C}}\subset\mathcal{A}$. Suppose there exists a potential function $\phi:\widetilde{\mathcal{C}}\to\mathbb{R}$ that satisfies $\phi(a)-\phi(a_{i}^{\prime},a_{-i})=\widetilde{u}_{i}(a)-\widetilde{u}_{i}(a_{i}^{\prime},a_{-i})$ for all $i\in[m]$, $a\in\widetilde{\mathcal{C}}$, and $a^{\prime}_{i}\in\widetilde{\mathcal{C}}_{i}(a_{-i})$. Then, the mixed strategy extension of this game with strategy space $\mathcal{X}\coloneqq\times_{i\in[m]}\Delta(\mathcal{A}_{i})$, utilities $u_{i}(x)\coloneqq\mathbb{E}_{a\sim x}[\widetilde{u}_{i}(a)]$, and constraints $c_{j}(x)\coloneqq\mathbb{E}_{a\sim x}[\widetilde{c}_{j}(a)]$ with thresholds $\alpha_{j}$ is a potential game with potential $\Phi(x)\coloneqq\mathbb{E}_{a\sim x}[\phi(a)]$. See Appendix F for a proof of this fact.
4.2 Independent learning protocol
We assume players to interact with the game for $T$ rounds. In each round $t$, for $0\leq t\leq T-1$, player $i$ chooses some strategy $\smash{x_{i}^{(t)}\in\mathcal{X}_{i}}$ and receives first-order feedback for the utility $\nabla_{x_{i}}u_{i}(x^{(t)})$, as well as zeroth- and first-order feedback for each constraint $j\in[b]$, that is, $c_{j}(x^{(t)})$ and $\nabla_{x_{i}}c_{j}(x^{(t)})$. Players do not observe other players’ strategies or utility feedback. Moreover, besides the game interaction, no communication or coordination (e.g., shared randomness) among players is allowed.
4.3 Interior point method
Access to a feasible initial strategy is crucial, as otherwise even a two-player game with bilinear constraints requires solving an NP-hard optimization problem (Witsenhausen,, 1986). For iterate $x^{(t)}\in\mathcal{C}$, we denote its feasibility margin by $\beta^{(t)}\coloneqq\min_{j\in[b]}\{c_{j}(x^{(t)})-\alpha_{j}\}$.
@@ 4.3 Interior point method
The initial strategy $x^{(0)}\in\mathcal{X}$ is strictly feasible, i.e., $\beta^{(0)}>0$, and for all $i\in[m]$, $x^{(0)}_{i}$ is initially known to player $i$.
To ensure feasibility of the iterates during the algorithm, we use log barrier regularized gradient descent, an interior point approach recently studied for nonconvex problems with nonconvex constraints (Hinder and Ye,, 2024; Usmanova et al.,, 2024). Under the assumed potential structure, we show that independently updated iterates remain in the interior of the joint feasible region.
@@
where $\eta>0$ is a parameter controlling the regularization. Observe that for each constraint $j\in[b]$, the log term serves as a barrier that tends to -$\infty$ as $c_{j}(z)$ approaches the threshold $\alpha_{j}$. Then, we propose the following independent log barrier regularized gradient ascent: Suppose in iteration $t$ for $0\leq t\leq T-1$, each player $i$ simultaneously performs the projected regularized gradient update
@@
4.4 Convergence Result
In order to prove our convergence result, we adopt standard assumptions on smoothness and Lipschitz continuity of utility functions, see also Mertikopoulos et al., (2024), and extend these regularity conditions to the constraints.
Assumption 4.
There exist constants $L$ and $M$ such that the functions $u_{i}$ for $i\in[m]$ and $c_{j}$ for $j\in[b]$ are all differentiable, $L$-Lipschitz continuous, and $M$-smooth in the joint strategies $x\in\mathcal{X}$.
Despite smoothness of utilities and constraints, $B^{\eta}_{i}(x)$ is not globally smooth due to unbounded growth of the log term towards the boundary of $\mathcal{C}$, which represents a key challenge in the analysis. However, we show that $B^{\eta}_{i}(x)$ is smooth over the joint strategy trajectory of independent log barrier regularized gradient ascent when using suitable adaptive stepsizes $\smash{\gamma^{(t)}}$.
Choice of stepsizes
Stepsizes are chosen as
where $M^{\eta,\beta^{(t)}}_{\Phi}\coloneqq M\sqrt{m}+\eta Mb(\beta^{(t)})^{-1}+\eta L^{2}b(\beta^{(t)})^{-2}$. The first term in (2) ensures that iterates remain within the feasible region. The second term bounds $\gamma^{(t)}$ by the inverse of the local smoothness parameter $M^{\eta,\beta^{(t)}}_{\Phi}$. We further describe these terms in the proof overview of Theorem 2 below.
@@ 4.4 Convergence Result
Let $\rho>0$, and let $\mathcal{J}_{\rho}(x)\coloneqq\{j\in[b]\mid c_{j}(x)-\alpha_{j}\leq\rho\}$ be the set of $\rho$-approximately active constraints for strategy $x\in\mathcal{C}$. There exists $\ell>0$, such that for all $x\in\mathcal{X}$ and all $j\in\mathcal{J}_{\rho}(x)$, there exists $\delta$ with $x+\delta\in\mathcal{X}$, $\lVert\delta\rVert\leq 1$, such that $\langle\delta,\nabla c_{j}(x)\rangle>\ell$.
@@
As the solution concept of our algorithms, we introduce an approximate notion of constrained Nash equilibrium. Note that the case $\epsilon=0$ recovers the exact notion from Section 2.
Definition 2.
For $\epsilon>0$, $x\in\mathcal{X}$ is a constrained $\epsilon$-approximate Nash equilibrium if $x\in\mathcal{C}$ and for all $i\in[m]$ and $x^{\prime}_{i}\in\mathcal{C}_{i}(x_{-i})$, $u_{i}(x^{\prime}_{i},x_{-i})\leq u_{i}(x)+\epsilon$.
Now, we are ready to state our main theorem about convergence to a constrained Nash equilibrium. {restatable}theoremthmpotential Let $\epsilon>0$ and set $\eta=\epsilon$. Let Assumptions 1, 3, 4, and 5 hold with $\rho\geq\eta$. For stepsizes $\{\gamma^{(t)}\}_{t=0}^{T-1}$ chosen as in (2), suppose we perform $T=\mathcal{O}(\epsilon^{-3})$ rounds of simultaneous updates as in (1). Then, there exists $0\leq t\leq T$ such that $x^{(t)}$ is a constrained $\epsilon$-approximate Nash equilibrium. We first discuss how this result relates to the existing literature. Then we provide an overview of our main ideas and the challenges of our analysis; for the full proof we refer to Appendix F.
Remarks.
@@
    The work of Anagnostides et al., (2022) also considers concave continuous action potential games, but without coupling constraints, establishing convergence to a Nash equilibrium at an $\mathcal{O}(\epsilon^{-2})$ best-iterate rate (see their Theorem B.6). Thus, handling coupling constraints via log barrier regularization comes at the cost of an additional $\mathcal{O}(\epsilon^{-1})$ factor. This is due to small adaptive stepsizes as iterates approach the boundary of the feasible region.
@@
    While we are not aware of any results on learning Nash equilibria in continuous action potential games with coupling constraints, a related result exists for constrained Markov potential games under an independent learning protocol similar to ours (Jordan et al.,, 2024). Unlike our approach, this result relies on a double-loop proximal-point-like algorithm that requires a so-called uniform Slater’s condition. Intermediate iterates are not guaranteed to remain within the feasible region. In the exact-gradient setting, a best-iterate convergence to a constrained $\epsilon$-approximate Nash equilibrium at a slower $\mathcal{O}(\epsilon^{-4})$ rate is proven.
@@
The main challenge of the analysis is to ensure feasibility of all iterates and smoothness of the objective along the trajectory despite the players’ independent updates. Towards this end, we observe that due to the cooperative structure of the log barrier regularization, $\Phi^{\eta}(x)\coloneqq\Phi(x)+\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$ is a potential function for the regularized game. Therefore, partial gradients are aligned with the potential gradient, i.e., $\nabla_{x_{i}}B^{\eta}_{i}(x)=\nabla_{x_{i}}\Phi^{\eta}(x)$ for all $i\in[m]$ which allows us to equivalently analyze our independent updates as a centralized log barrier regularized gradient ascent on $\Phi^{\eta}$.
Next, we reason about feasibility of the joint strategy iterates and smoothness of $\Phi^{\eta}$ along the gradient ascent trajectory. For $\beta^{(t)}$-feasible $x^{(t)}$, we give a Hessian bound $\lVert\nabla^{2}\Phi^{\eta}(x)\rVert\leq M^{\eta,\beta^{(t)}}_{\Phi}$. Moreover, we show that if an iterate $x^{(t)}$ is $\beta^{(t)}$-feasible, and stepsizes are chosen as in (2), then all strategies on the line segment $x^{(t)},x^{(t+1)}]$ are $(\beta^{(t)}/2)$-feasible. Consequently, $\Phi^{\eta}$ is $(M_{\Phi}^{\eta,\beta^{(t)}/2})$-smooth over $x^{(t)},x^{(t+1)}]$. Therefore, if we additionally ensure $\gamma^{(t)}\leq 1/(M_{\Phi}^{\eta,\beta^{(t)}/2})$, as in (2), a standard inequality from nonconvex smooth optimization (Bubeck et al., (2015), Lemma 3.6) shows sufficient increase proportional to $\gamma^{(t)}$. Nevertheless, $M_{\Phi}^{\eta,\beta^{(t)}/2}$ may grow arbitrarily and lead to diminishing stepsizes. Leveraging Assumption 5, we give a lower bound on $\gamma^{(t)}$ based on the fact that the log barrier term dominates and steers iterates away from the boundary of $\mathcal{C}$ whenever the feasibility margin $\beta^{(t)}$ is too small.
As a result, for our choices of $\eta$ and $T$, we obtain best-iterate convergence at an $\mathcal{O}(\epsilon^{-3})$ rate towards an $\epsilon$-stationary point $x^{\star}$ of $\Phi^{\eta}$. This implies a set of $\epsilon$-approximate KKT conditions to hold for the constrained problem $\max_{x\in\mathcal{C}}\Phi^{\eta}(x)$ at $x^{\star}$, which means $x^{\star}$ is an $\epsilon$-approximate constrained Nash equilibrium.
@@ 5 SIMULATIONS
To demonstrate the convergence of our method, we conduct numerical simulations on two continuous potential games with shared coupling constraints. The first is a simple two-player cooperative game played over a nonconvex feasible region. The second is a routing problem with link capacity constraints. We measure convergence in terms of $\text{Nash-Gap}_{i}(x)\coloneqq\max_{x_{i}^{\prime}\in\mathcal{C}_{i}(x_{-i})}u_{i}(x_{i}^{\prime},x_{-i})-u_{i}(x)$ for each player $i$. This can be computed by solving the single-player constrained optimization problem where $x_{-i}$ remains fixed. We further define $\text{Nash-Gap}(x)\coloneqq\max_{i\in[m]}\text{Nash-Gap}_{i}(x)$. This serves as a measure of convergence, as for any $\epsilon>0$, if $\text{Nash-Gap}(x)\coloneqq\max_{i\in[m]}\text{Nash-Gap}_{i}(x)\leq\epsilon$, then $x$ is a constrained $\epsilon$-approximate Nash equilibrium.
Cooperative game
@@
Figure 3: Strategy trajectory in the utility landscape of the cooperative game; colors indicate the log barrier value; grey is infeasible.
Starting from an initial feasible point, we apply our log barrier regularized independent gradient ascent to the constrained potential game. Figure 3 visualizes the joint strategy trajectory within the utility landscape where colors indicate the log barrier value. The trajectory (in red) remains within the nonconvex feasible region throughout the iterations, aligning with our theoretical observations. In Figure 4, we plot the $\text{Nash-Gap}_{i}(x)$ for each player $i\in[1,2]$ and observe convergence to a strategy near the boundary of $\mathcal{C}$ that forms an approximate constrained Nash equilibrium.
Figure 4: Nash gaps and constraint values over iterations $t$ for the cooperative game.
@@
Beyond the cooperative case, we turn to a classical distributed routing problem, a canonical example of a potential game well studied in the literature on network routing (Menache and Ozdaglar,, 2011). We extend this model by incorporating coupled capacity constraints. Specifically, 5 players, one corresponding to each color in Figure 5, route some amount of flow through 5 paths in the network, each associated to the respective player ($R_{1}$: blue, $R_{2}$: red, $R_{3}$: brown, $R_{4}$: orange, $R_{5}$: green). For $i\in[5]$, we let $x_{i}\in[0,10]$ denote the flow on player $i$’s path $R_{i}$.
@@
Starting from feasible flows $x_{i}=\frac{1}{2}$ for all $i\in[5]$, Figure 6 shows the convergence of log barrier regularized independent gradient ascent to a constrained Nash equilibrium of the routing game. We point out that, again, all iterates remain within the feasible region.